# Towards Reliable Forgetting: A Survey on Machine Unlearning Verification, Challenges, and Future Directions

### v1

@@ 3.2. Privacy Analysis
Privacy analysis verification methods implement unlearning verification by analyzing whether the unlearned model retains any privacy information about the forgotten data $D_{f}$. In this method, the query set ${Q}_{D_{f}}$ consists of data points from $D_{f}$ itself. The behavior $\mathcal{B}$ refers to whether the model retains any privacy information about $D_{f}$. The fundamental principle behind constructing the reference behavior set ${T}_{\neg D_{f}}$ is that models that have never encountered $D_{f}$ data should not leak any membership information about $D_{f}$ under the privacy analysis module. If the unlearned model leaks privacy information about $D_{f}$, it indicates that the model has failed the privacy analysis verification. This type of verification primarily utilizes the following privacy analysis techniques: data inference, data reconstruction, and weight analysis.
@@ 3.2.1. Data Inference
Existing data inference approaches for unlearning verification can be broadly categorized into two types: Membership Inference (MI) and Non-Membership Inference (NMI). MI methods aim to determine whether the target model has encountered a specific data point during training, and have been widely used to distinguish between training and unseen data (Chen et al., 2021; Shokri et al., 2017). Some studies further adapt MI to the unlearning verification setting, where a data point is considered successfully unlearned if it is inferred as a non-member after the unlearning request (Tu et al., 2024; Di et al., 2024; Sharma et al., 2024; Liu et al., 2023b; Seo et al., 2025). Additionally, Kim et al. (2025) proposes an implicit membership inference framework tailored for large vision-language models in text generation tasks, aimed at evaluating whether the model has truly forgotten sensitive information related to specific entities. In this approach, Wikipedia-style summaries of target entities are used as proxies for private information. By measuring the semantic similarity between the model’s generated output and the entity summary, the method indirectly assesses whether memorization persists. Specifically, if the output shows significantly higher similarity to the target entity’s description than to that of unrelated entities, it is considered indicative of residual membership bias, suggesting incomplete forgetting. This method does not require access to model parameters and is applicable in black-box settings.
In contrast, NMI focuses on identifying data points that were previously memorized by the model but should have been forgotten following an unlearning operation. Some works (Chen et al., 2021; Hayes et al., 2024) have proposed inference methods that leverage discrepancies between the outputs of the original and unlearned models. Lu et al. (Lu et al., 2022) extend this idea to label-only black-box settings, which infer unlearning success based on the amount of perturbation required to change predicted labels, even when only hard labels are accessible. Further, Wang et al. (2024a) systematically introduce NMI into the unlearning verification framework, utilizing model state transitions to enhance the reliability of unlearning validation. To enable more efficient verification, their method operates in a white-box setting and directly performs inference on the target model, eliminating the need to train multiple shadow models.
@@ 3.2.2. Data Reconstruction
Data reconstruction is a class of privacy analysis techniques that aim to recover specific features of target data from a given model (Fredrikson et al., 2015; Buzaglo et al., 2023; Haim et al., 2022). Some studies (Graves et al., 2021; Xu et al., 2024a) show that such techniques can be used to evaluate whether a model successfully forgets particular data. Graves et al. (Graves et al., 2021) demonstrate that class-level reconstruction reveals whether a model retains the semantic information of certain categories; however, it lacks the granularity needed to verify forgetting at the sample level. To address this limitation, Xu et al. (Xu et al., 2024a) propose a parameter-driven reconstruction framework that formulates an optimization objective based on implicit bias theory. Their method enables direct recovery of previously seen training samples from model parameters, without requiring labels or external triggers, making it well-suited for general-purpose unlearning verification.
The core idea behind these approaches is that a properly unlearned model should no longer be able to reconstruct information associated with the forgotten data. If such information remains recoverable, it indicates that the unlearning process is incomplete. Recently, some works (Bertran et al., 2024; Hu et al., 2024) have introduced a more direct verification approach: attempting to reconstruct the specific data samples that have been requested for deletion. If a reconstructed sample matches one of these forgotten targets, the data is considered not effectively forgotten. Bertran et al. (2024) implement reconstruction-based verification for unlearned samples in simple models; however, their approach does not scale well to complex deep neural networks. To address this limitation, Hu et al. (2024) apply gradient inversion techniques to enable the reconstruction of unlearned samples in more complex architectures, demonstrating improved effectiveness in realistic settings.
@@ 3.3. Model Performance
Model performance unlearning verification methods aim to evaluate the effectiveness of unlearning by assessing the model’s performance on forgotten data. In this method, the query set ${Q}_{D_{f}}$ consists of data points from $D_{f}$. The behavior $\mathcal{B}$ represents the model’s performance on the query set ${Q}_{D_{f}}$. The intuition behind constructing the behavioral set ${T}_{\neg D_{f}}$ is that models exposed to $D_{f}$ exhibit significantly different performance on tasks related to $D_{f}$ compared to those that have not encountered it. Existing performance-based metrics for unlearning verification include task quality, relearn time, model output, and model sensitivity.
@@
Task quality refers to a model’s ability to perform its intended function on a given dataset. This is typically measured by performance metrics specific to the task, such as classification accuracy for supervised learning, semantic coherence for text generation, or perceptual quality for image synthesis. Task quality evaluation on the forgotten dataset is a widely used and intuitive approach for assessing unlearning effectiveness (Graves et al., 2021; Mehta et al., 2022; Chundawat et al., 2023a, b; Kurmanji et al., 2023). The underlying idea is that if the model’s task quality on the forgotten data significantly decreases, this indicates that the model has reduced or lost its reliance on those samples, suggesting that forgetting has occurred. In addition, Nguyen et al. (2022b) propose comparing the task quality of the unlearned model with that of a model retrained from scratch on the retained dataset, which has never seen the erased data. Although this method provides a useful benchmark, it also introduces additional computational overhead due to the need for retraining. Recently, Kim et al. (2025) evaluates forgetting by measuring task quality in text generation. It assesses the semantic similarity between the model’s output and the target entity’s description. If the output closely matches the target while diverging from non-targets, it indicates that the model retains knowledge and has not effectively forgotten.
@@
Relearn time refers to the number of training epochs required for a model to relearn and restore its performance on forgotten data. Studies (Fan et al., 2023; Golatkar et al., 2020b, a; Chundawat et al., 2023b) use a loss function to evaluate the model’s performance on the target data and set a threshold below which the model’s loss on the forgotten data points is considered sufficiently low. When the model’s loss falls below this threshold, it is deemed to have successfully relearned the forgotten data. A higher number of training epochs generally indicates better forgetting performance by the model.
@@
The performance of the model is largely reflected in its output. Some studies attempt to demonstrate the model’s forgetting effect by measuring the difference in outputs between the model before and after forgetting. For example, some studies (Golatkar et al., 2020b, a) use the distance between the outputs of two models for forgotten data or forgotten data features. Peste et al. (2021) propose measuring unlearning effectiveness by calculating the $L_{1}$ distance between the confusion matrices of the original and unlearned models, and Wang et al. (2025b) audit unlearning by analyzing the posterior difference between the model outputs before and after unlearning on the same data. Jeon et al. (2024b) quantify the residual forgotten information retained by the model by measuring the mutual information between the intermediate layer outputs and the labels of the data to be forgotten. Recently, Vidal et al. (2024) propose an unlearning verification method that leverages explainability tools to assess whether a model retains memory of target information. By analyzing differences in the model’s responses before and after unlearning, the method uses attribution techniques to generate heatmaps based on intermediate activations and model outputs, thereby quantifying changes in the model’s attention to key input regions.
@@
Model sensitivity refers to the degree to which a model’s output behavior or parameter gradients change when exposed to a given input, thereby reflecting the model’s dependency on that input. This concept is concretely modeled in some works to enable unlearning verification. Becker and Liebig (2022) propose a sensitivity-based evaluation method grounded in Fisher information to assess the effectiveness of machine unlearning. Specifically, the trace of the Fisher Information Matrix is used as a core metric to quantify the average gradient response of model parameters to the target data. A higher sensitivity indicates that the model has stably encoded the target data, whereas a successful unlearning process should reduce the Fisher information, leading to decreased sensitivity and increased epistemic uncertainty. This method assumes a white-box setting, as it requires access to gradient information, and supports efficient estimation through gradient-based approximations and computable upper bounds. In contrast, Zhou et al. (2025a) propose the TruVRF framework, which adopts an alternative definition of model sensitivity by measuring the magnitude of parameter updates during fine-tuning. The core idea is to infer whether the model retains memory of the target data based on how much its parameters need to adjust. If the model has not seen the data before, or has successfully forgotten it, fine-tuning will induce substantial parameter changes, resulting in high sensitivity. In contrast, if the model remembers the data, only minimal updates are required, corresponding to low sensitivity. This method also assumes a white-box setting, as it relies on accessing parameter-level changes during fine-tuning. Although these methods differ in how they define and interpret sensitivity, they share a common principle: modeling the response strength of model parameters to input data as a proxy for memory retention. It is important to note that all such approaches inherently assume white-box access due to their reliance on internal parameter visibility, which may limit their applicability in certain real-world scenarios.
@@ 4. Parametric Unlearning Verification
Core Dependency Based on the indistinguishability of model parameters under neighboring datasets, formalized via differential privacy. Based on the ability to reconstruct the training trajectory from recorded updates and batch information.
Table 6. Key Symbols Newly Introduced for Parametric Unlearning Verification
@@
$N(x,y)$ Nearest retained sample to $(x,y)$ from the same class
$x$ A single data point to be removed from the dataset
@@
$w_{t}$ Model parameters at training step $t$
@@
$w_{t}$ Model parameters at step $t$
$(w_{i},x_{i},g_{i})$ Training log tuple.
@@
4.1. Differential Privacy Verification.
@@
Inspired by the concept of differential privacy, studies (Guo et al., 2019; Ginart et al., 2019) introduce a probabilistic formalization that limits the difference between the model weights returned by the machine unlearning and retraining algorithms, providing provable guarantees of forgetting for the removed data. The following definitions are proposed:
Definition 0 ($\epsilon$-Certified Forgetting (Guo et al., 2019)).
Let $A:D\to H$ be a randomized learning algorithm, where the training dataset $D$ induces a probability distribution over the hypothesis space $H$. Define a data removal mechanism $M$, which aims to remove the influence of a sample $x\in D$ from the output of $A(D)$. If, for all possible training datasets $D\subseteq X$, all possible removal samples $x\in D$, and all possible hypothesis sets $S\subseteq H$, the following inequality holds:
$e^{-\epsilon}\leq\frac{P(M(A(D),D,x)\in S)}{P(A(D\setminus x)\in S)}\leq e^{\epsilon}$
@@
This property ensures that the model distribution after removing $x$ is close to the model distribution that has never seen $x$, making it difficult for an adversary to distinguish whether $x$ was used during training. Guo et al. (Guo et al., 2019) further relax the conditions and introduced the following result:
Definition 0 ($(\epsilon,\delta)$-Certified Forgetting (Guo et al., 2019)).
For $\delta>0$, if the data removal mechanism $M$ satisfies the following relaxed conditions:
$P(M(A(D),D,x)\in S)\leq e^{\epsilon}P(A(D\setminus x)\in S)+\delta,$
$P(A(D\setminus x)\in S)\leq e^{\epsilon}P(M(A(D),D,x)\in S)+\delta.$
then $M$ is said to achieve $(\epsilon,\delta)$-Certified Forgetting.
Here, $\delta$ controls the upper bound on the probability of failure, allowing for a small divergence probability, thereby relaxing the strict constraints of $\epsilon$-Certified Forgetting.
It can be observed that achieving certified forgetting inherently enables verification of the unlearned model in the parameter space. We refer to this line of work as differential privacy unlearning verification. Several methods have been proposed in this context, which can be broadly categorized into first-order approaches based on function gradients and second-order approaches that leverage Hessian information or its approximations, as outlined below.
@@
First-order methods rely on gradient-related information to achieve provable unlearning. For example, Chien et al. (2024b) combine differential privacy and proposes an approximate unlearning framework based on projection noise gradient descent. This framework is suitable for empirical risk minimization problems with smooth objectives, providing provable forgetting capabilities. Chien et al. (2024a) propose an approximate unlearning method based on projected noisy stochastic gradient descent, offering certified unlearning guarantees by tracking the Wasserstein distance of the parameter distribution. This approach extends to multiple unlearning requests, including sequential and batch settings, and achieves better loss bounds through tighter analysis of the $W_{\infty}$ distance. Additionally, Wang and Wang (2025) introduce provable forgetting in federated learning scenarios by adding Gaussian noise to the gradient during both model training and unlearning, and applying a specific gradient ascent algorithm during the unlearning update. Huynh et al. (2025) propose the CFRU algorithm for provable forgetting in federated recommendation systems. This method rolls back and removes the target client’s historical updates, uses a sampling strategy to reduce storage requirements, estimates the deletion bias using Lipschitz conditions, and compensates for model errors through an iterative scheme without the need for additional training. Wei et al. (2024) provide the first theoretical guarantee for learning in pre-training and fine-tuning paradigms, focusing on bag-of-words language models, which can be applied to downstream tasks such as retrieval and classification.
The above methods focus on convex or strongly convex functions. Recently, Mu and Klabjan (2024) propose a provable unlearning method for non-convex functions. This method performs unlearning by ”rewinding” to earlier steps in the learning process before applying gradient descent on the loss function for retained data points, achieving a better balance between privacy, utility, and computational complexity. The method is shown to provide generalization guarantees for non-convex functions that satisfy the Polyak-Lojasiewicz inequality.
@@
These methods rely on Hessian or approximations of Hessian-related information to provide certified unlearning. Guo et al. (Guo et al., 2019) first introduce a provable deletion mechanism for $L_{2}$-regularized linear models, applicable to models trained with various convex loss functions (e.g., logistic regression). This mechanism significantly reduces the impact of deleted data points by applying a Newton step to the model parameters, causing the residual error to decrease quadratically with the size of the training set. To better prove forgetting, they further introduce a method based on random perturbation of the training loss to mask these residual errors. This method requires retraining the model, and to reduce computational costs, Sekhari et al. (2021) propose an algorithm that enables sample deletion without accessing the original training data. The algorithm estimates the impact of deleted data on the model using the Hessian matrix and adjusts through noise perturbation, ensuring provable forgetting. Recently, Jin et al. (2023) introduce a certified unlearning strategy for federated learning. This strategy performs unlearning using Newton updates on the server side, simplifying the training process with a linear loss function, requiring no client participation or additional storage. This method effectively controls the difference between the unlearned weights and the original optimal weights, enabling efficient and certified unlearning.
@@
Although the above methods provide proofs for forgetting, they are not practical. They only consider convexity assumptions, involve costly Hessian matrix computations, and assume the learning model is in the empirical risk minimization setting. To enhance practicality in deep learning scenarios, Zhang et al. (2024b) investigate the problem of certified forgetting without relying on convexity assumptions. The study shows that by making simple modifications to Newton updates, an approximation error bound for non-convex functions can be established, and various approximation unlearning strategies for convex models can be adapted to improve their robustness in deep neural networks. Additionally, to improve efficiency, they propose a computationally efficient method to estimate the inverse Hessian in Newton updates and prove that this method retains bounded approximation errors. Qiao et al. (2024) introduce an innovative Hessian-free second-order unlearning method designed to overcome the limitations of traditional second-order algorithms, such as high Hessian computation costs, inability to handle multiple deletion requests online, and reliance on convexity assumptions. By analyzing each sample’s impact on the training trajectory, this method avoids explicitly computing the Hessian matrix, enabling efficient online deletion request handling. Compared to existing methods, this approach does not require the learning model to be an empirical risk minimizer and offers better scalability. Furthermore, Liu et al. (2023a) develop a new certified machine unlearning algorithm for min-max models, achieving min-max unlearning steps with fully Newton-based updates using Hessian and incorporating Gaussian noise through differential privacy. To ensure certified unlearning, this method injects calibrated Gaussian noise by carefully analyzing the similarity between min-max unlearning variables and retraining variables, applicable to (strongly-)convex-(strongly-)concave loss functions.

### v2

@@ 3.2. Privacy Analysis
Privacy analysis verification evaluates the effectiveness of unlearning by constructing a privacy analysis module that examines the leakage of sensitive information related to $D_{f}$. In this approach, the query set $Q_{D_{f}}$ typically consists of data points sampled directly from $D_{f}$, and the behavioral function $\mathcal{B}$ measures the model’s responses under the privacy analysis module, reflecting its potential retention of information about $D_{f}$. The core intuition is that a model which has not forgotten $D_{f}$ and a model $\mathbf{w}^{*}$ that has never seen $D_{f}$ should exhibit significantly different privacy signals when evaluated through the privacy analysis module on $Q_{D_{f}}$. For example, when using membership inference for verification, one can infer that $D_{f}$ belongs to the training set from the non-unlearned model, whereas $D_{f}$ would be inferred as a non-member in the model $\mathbf{w}^{*}$. This type of verification primarily utilizes the following privacy analysis techniques: data inference and and data reconstruction.
@@ 3.2.1. Data Inference
Existing data inference approaches for unlearning verification can be broadly categorized into two types: Membership Inference (MI) and Non-Membership Inference (NMI). MI methods aim to determine whether the target model has encountered a specific data point during training, and have been widely used to distinguish between training and unseen data (Chen et al., 2021; Shokri et al., 2017). Some studies further adapted MI to the unlearning verification setting, where a data point is considered successfully unlearned if it is inferred as a non-member after the unlearning request (Tu et al., 2024; Di et al., 2025; Sharma et al., 2024; Liu et al., 2023b; Seo et al., 2025). Additionally, Kim et al. (2025) proposed an implicit membership inference framework tailored for large vision-language models in text generation tasks, aimed at evaluating whether the model has truly forgotten sensitive information related to specific entities. In this approach, Wikipedia-style summaries of target entities are used as proxies for private information. By measuring the semantic similarity between the model’s generated output and the entity summary, the method indirectly assesses whether memorization persists. Specifically, if the output shows significantly higher similarity to the target entity’s description than to that of unrelated entities, it is considered indicative of residual membership bias, suggesting incomplete forgetting. This method does not require access to model parameters and is applicable in black-box settings.
MI focuses on identifying which data the unlearned model still retains, relying solely on access to the unlearned model. In contrast, NMI aims to identify which data points have been forgotten by comparing model responses before and after unlearning, thus requiring access to both the original and unlearned models. Some works (Chen et al., 2021; Hayes et al., 2024) have proposed inference methods that leverage discrepancies between the outputs of the original and unlearned models. Lu et al. (Lu et al., 2022) extended this idea to label-only black-box settings, which infer unlearning success based on the amount of perturbation required to change predicted labels, even when only hard labels are accessible. Further, Wang et al. (2024a) systematically introduced NMI into the unlearning verification framework, utilizing model state transitions to enhance the reliability of unlearning validation. To enable more efficient verification, their method operates in a white-box setting and directly performs inference on the target model, eliminating the need to train multiple shadow models.
@@ 3.2.2. Data Reconstruction
Data reconstruction is a class of privacy analysis techniques that aim to recover specific features of target data from a given model (Fredrikson et al., 2015; Buzaglo et al., 2023; Haim et al., 2022). Some studies (Graves et al., 2021; Xu et al., 2024a) show that such techniques can be used to evaluate whether a model successfully forgets particular data. Graves et al. (Graves et al., 2021) demonstrated that class-level reconstruction reveals whether a model retains the semantic information of certain categories; however, it lacks the granularity needed to verify forgetting at the sample level. To address this limitation, Xu et al. (Xu et al., 2024a) proposed a parameter-driven reconstruction framework that formulates an optimization objective based on implicit bias theory. Their method enables direct recovery of previously seen training samples from model parameters, without requiring labels or external triggers, making it well-suited for general-purpose unlearning verification.
The core idea behind these approaches is that a properly unlearned model should no longer be able to reconstruct information associated with the forgotten data. If such information remains recoverable, it indicates that the unlearning process is incomplete. Recently, some works (Bertran et al., 2024; Hu et al., 2024) have introduced a more direct verification approach: attempting to reconstruct the specific data samples that have been requested for deletion. If a reconstructed sample matches one of these forgotten targets, the data is considered not effectively forgotten. Bertran et al. (2024) implemented reconstruction-based verification for unlearned samples in simple models; however, their approach does not scale well to complex deep neural networks. To address this limitation, Hu et al. (2024) applied gradient inversion techniques to enable the reconstruction of unlearned samples in more complex architectures, demonstrating improved effectiveness in realistic settings.
@@ 3.3. Model Performance
Model performance unlearning verification methods aim to evaluate the effectiveness of unlearning by assessing the model’s performance on forgotten data. In this method, the query set ${Q}_{D_{f}}$ consists of data points from $D_{f}$. The behavior $\mathcal{B}$ represents the model’s performance on the query set ${Q}_{D_{f}}$. The core intuition behind such methods is that models exposed to $D_{f}$ will exhibit different performance compared to models that have never seen $D_{f}$. Existing performance-based unlearning verification metrics include task quality, retraining time, model outputs, and model sensitivity.
@@
Task quality refers to a model’s ability to perform its intended function on a given dataset. This is typically measured by performance metrics specific to the task, such as classification accuracy for supervised learning, semantic coherence for text generation, or perceptual quality for image synthesis. Task quality evaluation on the forgotten dataset is a widely used and intuitive approach for assessing unlearning effectiveness (Graves et al., 2021; Mehta et al., 2022; Chundawat et al., 2023a, b; Kurmanji et al., 2023). The underlying idea is that if the model’s task quality on the forgotten data significantly decreases, this indicates that the model has reduced or lost its reliance on those samples, suggesting that forgetting has occurred. In addition, Nguyen et al. (2022b) proposed comparing the task quality of the unlearned model with that of a model retrained from scratch on the retained dataset, which has never seen the erased data. Although this method provides a useful benchmark, it also introduces additional computational overhead due to the need for retraining. Recently, Kim et al. (2025) evaluated forgetting by measuring task quality in text generation. It assesses the semantic similarity between the model’s output and the target entity’s description. If the output closely matches the target while diverging from non-targets, it indicates that the model retains knowledge and has not effectively forgotten.
@@
Relearn time refers to the number of training epochs required for a model to relearn and restore its performance on forgotten data. Studies (Fan et al., 2024; Golatkar et al., 2020b, a; Chundawat et al., 2023b) use a loss function to evaluate the model’s performance on the target data and set a threshold below which the model’s loss on the forgotten data points is considered sufficiently low. When the model’s loss falls below this threshold, it is deemed to have successfully relearned the forgotten data. A higher number of training epochs generally indicates better forgetting performance by the model.
@@
The performance of the model is largely reflected in its output. Some studies attempt to demonstrate the model’s forgetting effect by measuring the difference in outputs between the model before and after forgetting. For example, some studies (Golatkar et al., 2020b, a; Chundawat et al., 2023a; Nguyen et al., 2020) use the distance between the outputs of two models for forgotten data or forgotten data features. Peste et al. (2021) proposed measuring unlearning effectiveness by calculating the $L_{1}$ distance between the confusion matrices of the original and unlearned models, and Wang et al. (2025b) audited unlearning by analyzing the posterior difference between the model outputs before and after unlearning on the same data. Jeon et al. (2024) quantified the residual forgotten information retained by the model by measuring the mutual information between the intermediate layer outputs and the labels of the data to be forgotten. Recently, Vidal et al. (2024) proposed an unlearning verification method that leverages explainability tools to assess whether a model retains memory of target information. By analyzing differences in the model’s responses before and after unlearning, the method uses attribution techniques to generate heatmaps based on intermediate activations and model outputs, thereby quantifying changes in the model’s attention to key input regions.
@@
Model sensitivity refers to the degree to which a model’s output behavior or parameter gradients change when exposed to a given input, thereby reflecting the model’s dependency on that input. This concept is concretely modeled in some works to enable unlearning verification. Becker and Liebig (2022) proposed a sensitivity-based evaluation method grounded in Fisher information to assess the effectiveness of machine unlearning. Specifically, the trace of the Fisher Information Matrix is used as a core metric to quantify the average gradient response of model parameters to the target data. A higher sensitivity indicates that the model has stably encoded the target data, whereas a successful unlearning process should reduce the Fisher information, leading to decreased sensitivity and increased epistemic uncertainty. This method assumes a white-box setting, as it requires access to gradient information, and supports efficient estimation through gradient-based approximations and computable upper bounds. In contrast, Zhou et al. (2025a) proposed the TruVRF framework, which adopts an alternative definition of model sensitivity by measuring the magnitude of parameter updates during fine-tuning. The core idea is to infer whether the model retains memory of the target data based on how much its parameters need to adjust. If the model has not seen the data before, or has successfully forgotten it, fine-tuning will induce substantial parameter changes, resulting in high sensitivity. In contrast, if the model remembers the data, only minimal updates are required, corresponding to low sensitivity. This method also assumes a white-box setting, as it relies on accessing parameter-level changes during fine-tuning. Although these methods differ in how they define and interpret sensitivity, they share a common principle: modeling the response strength of model parameters to input data as a proxy for memory retention. It is important to note that all such approaches inherently assume white-box access due to their reliance on internal parameter visibility, which may limit their applicability in certain real-world scenarios.
@@ 4. Parametric Unlearning Verification
Table 5. Key symbols newly introduced for parametric unlearning verification.
@@
$N(\mathbf{x},y)$ Nearest retained sample to $(\mathbf{x},y)$ from the same class
$\mathbf{x}$ A single data point to be removed from the dataset
@@
$\mathbf{w}_{t}$ Model parameters at training step $t$
@@
$(\mathbf{w}_{i},\mathbf{x}_{i},g_{i})$ Training log tuple
@@
4.1. Differential Privacy Verification
@@
Inspired by the concept of differential privacy, studies (Guo et al., 2020; Ginart et al., 2019) introduce a probabilistic formalization that limits the difference between the model weights returned by the machine unlearning and retraining algorithms, providing provable guarantees of forgetting for the removed data. The following definitions are proposed:
Definition 0 ($\epsilon$-Certified Forgetting (Guo et al., 2020)).
Let $A:D\to H$ be a randomized learning algorithm, where the training dataset $D$ induces a probability distribution over the hypothesis space $H$. Define a data removal mechanism $M$, which aims to remove the influence of a sample $\mathbf{x}\in D$ from the output of $A(D)$. If, for all possible training datasets $D\subseteq X$, all possible removal samples $\mathbf{x}\in D$, and all possible hypothesis sets $S\subseteq H$, the following inequality holds:
$e^{-\epsilon}\leq\frac{P(M(A(D),D,\mathbf{x})\in S)}{P(A(D\setminus\mathbf{x})\in S)}\leq e^{\epsilon}$
@@
This property ensures that the model distribution after removing $\mathbf{x}$ is close to the model distribution that has never seen $\mathbf{x}$, making it difficult for an adversary to distinguish whether $\mathbf{x}$ was used during training. Guo et al. (Guo et al., 2020) further relax the conditions and introduced the following result:
Definition 0 ($(\epsilon,\delta)$-Certified Forgetting (Guo et al., 2020)).
For $\delta>0$, if the data removal mechanism $M$ satisfies the following relaxed conditions: {IEEEeqnarray}rCl & P(M(A(D), D, x) S) e^ϵ P(A(D x) S) + δ,
P(A(D
x) S) e^ϵ P(M(A(D), D, x) S) + δ, then $M$ is said to achieve $(\epsilon,\delta)$-Certified Forgetting.
Here, $\delta$ controls the upper bound on the probability of failure, allowing for a small divergence probability, thereby relaxing the strict constraints of $\epsilon$-Certified Forgetting. The above definition concerns differential privacy guarantees for individual data points. To generalize certified forgetting to the multi-point deletion setting, several works (Guo et al., 2020; Zhang et al., 2024b) have proposed treating the set of data to be removed as a deletion unit. Building on this idea, subsequent studies (Wang and Wang, 2025; Huynh et al., 2025) have extended certified unlearning to the federated learning setting by enabling client-level unlearning, where a client’s entire dataset is treated as a single unit of deletion in the unlearning process.
It can be observed that this class of methods for achieving certified forgetting inherently has these verification capabilities within the model’s parameter space. We refer to this line of work as differential privacy unlearning verification. Several methods have been proposed in this context, which can be broadly categorized into first-order approaches based on function gradients and second-order approaches that leverage Hessian information or its approximations, as outlined below.
Core Dependency Based on the indistinguishability of model parameters under neighboring datasets, formalized via differential privacy. Based on the ability to reconstruct the training trajectory from recorded updates.
@@
First-order methods rely on gradient-related information to achieve provable unlearning. For example, Chien et al. (2024b) combined differential privacy and proposed an approximate unlearning framework based on projection noise gradient descent. This framework is suitable for empirical risk minimization problems with smooth objectives, providing provable forgetting capabilities. Chien et al. (2024a) proposed an approximate unlearning method based on projected noisy stochastic gradient descent, offering certified unlearning guarantees by tracking the Wasserstein distance of the parameter distribution. This approach extends to multiple unlearning requests, including sequential and batch settings, and achieves better loss bounds through tighter analysis of the $W_{\infty}$ distance. Additionally, Wang and Wang (2025) introduced provable forgetting in federated learning scenarios by adding Gaussian noise to the gradient during both model training and unlearning, and applying a specific gradient ascent algorithm during the unlearning update. Huynh et al. (2025) proposed the CFRU algorithm for provable forgetting in federated recommendation systems. This method rolls back and removes the target client’s historical updates, uses a sampling strategy to reduce storage requirements, estimates the deletion bias using Lipschitz conditions, and compensates for model errors through an iterative scheme without the need for additional training. Wei et al. (2024) provided the first theoretical guarantee for learning in pre-training and fine-tuning paradigms, focusing on bag-of-words language models, which can be applied to downstream tasks such as retrieval and classification.
The above methods focus on convex or strongly convex functions. Recently, Mu and Klabjan (2024) proposed a provable unlearning method for non-convex functions. This method performs unlearning by ”rewinding” to earlier steps in the learning process before applying gradient descent on the loss function for retained data points, achieving a better balance between privacy, utility, and computational complexity. The method is shown to provide generalization guarantees for non-convex functions that satisfy the Polyak-Lojasiewicz inequality.
@@
These methods rely on Hessian or approximations of Hessian-related information to provide certified unlearning. Guo et al. (Guo et al., 2020) first introduced a provable deletion mechanism for $L_{2}$-regularized linear models, applicable to models trained with various convex loss functions (e.g., logistic regression). This mechanism significantly reduces the impact of deleted data points by applying a Newton step to the model parameters, causing the residual error to decrease quadratically with the size of the training set. To better prove forgetting, they further introduce a method based on random perturbation of the training loss to mask these residual errors. This method requires retraining the model, and to reduce computational costs, Sekhari et al. (2021) proposed an algorithm that enables sample deletion without accessing the original training data. The algorithm estimates the impact of deleted data on the model using the Hessian matrix and adjusts through noise perturbation, ensuring provable forgetting. Recently, Jin et al. (2023) introduced a certified unlearning strategy for federated learning. This strategy performs unlearning using Newton updates on the server side, simplifying the training process with a linear loss function, requiring no client participation or additional storage. This method effectively controls the difference between the unlearned weights and the original optimal weights, enabling efficient and certified unlearning.
@@
Although the above methods provide proofs for forgetting, they are not practical. They only consider convexity assumptions, involve costly Hessian matrix computations, and assume the learning model is in the empirical risk minimization setting. To enhance practicality in deep learning scenarios, Zhang et al. (2024b) investigated the problem of certified forgetting without relying on convexity assumptions. The study shows that by making simple modifications to Newton updates, an approximation error bound for non-convex functions can be established, and various approximation unlearning strategies for convex models can be adapted to improve their robustness in deep neural networks. Additionally, to improve efficiency, they propose a computationally efficient method to estimate the inverse Hessian in Newton updates and prove that this method retains bounded approximation errors. Qiao et al. (2024) introduced an innovative Hessian-free second-order unlearning method designed to overcome the limitations of traditional second-order algorithms, such as high Hessian computation costs, inability to handle multiple deletion requests online, and reliance on convexity assumptions. By analyzing each sample’s impact on the training trajectory, this method avoids explicitly computing the Hessian matrix, enabling efficient online deletion request handling. Compared to existing methods, this approach does not require the learning model to be an empirical risk minimizer and offers better scalability. Furthermore, Liu et al. (2023a) developed a new certified machine unlearning algorithm for min-max models, achieving min-max unlearning steps with fully Newton-based updates using Hessian and incorporating Gaussian noise through differential privacy. To ensure certified unlearning, this method injects calibrated Gaussian noise by carefully analyzing the similarity between min-max unlearning variables and retraining variables, applicable to (strongly-)convex-(strongly-)concave loss functions.